\begin{thebibliography}{999}

\bibitem{CCLP92}
Chambre, P.L.; Chernoff, P.R.; Lehman, R.S.; Pinney, E.J.
\newblock University of California: In Memoriam, 1992 / Ren\'e J. De Vogelaere,
  Mathematics: Berkeley,  1992. 
\newblock  \url{http://texts.cdlib.org/view?docId=hb7c6007sj&chunk.id=div00013&brand=calisphere&doc.view=entire_text}

\bibitem{DeVo56}
{De Vogelaere}, R.
\newblock Methods of Integration Which Preserve the Contact Transformation
  Property of {H}amiltonian Equations.
\newblock Technical Report~4, Department of Mathematics, University of Notre
  Dame,  1956. \quad
\newblock \url{https://curate.nd.edu/downloads/6682x34921d}.

\bibitem{Ruth83}
Ruth, R.D.
\newblock A Canonical Integration Technique.
\newblock {\em IEEE Trans.\ Nucl.\ Sci.} {\bf 1983}, {\em NS-30},~2669--2671.

\bibitem{ChSc90}
Channell, P.J.; Scovel, J.C.
\newblock Symplectic Integration of \mbox{Hamiltonian} Systems.
\newblock {\em Nonlinearity} {\bf 1990}, {\em 3},~231--259.

\bibitem{Chan83}
Channell, P.
\newblock Symplectic integration algorithms.
\newblock Technical Report AT-6, ATN-83-9, Los Alamos National Laboratory,
  1983.

\bibitem{Feng85}
Feng, K.
\newblock On Difference Schemes and Symplectic Geometry.
\newblock  Proceedings of the 1984 Beijing Symposium on Differential Geometry
  and Differential Equations{\rm ---}Computation of Partial Differential
  Equations; Feng, K., Ed.; Science Press: Beijing,  1985; pp. 42--58.

\bibitem{Feng86}
Feng, K.
\newblock Difference Schemes for {H}amiltonian Formalism and Symplectic
  Geometry.
\newblock {\em J. Comput.\ Math.} {\bf 1986}, {\em 4},~279--289.

\bibitem{GeMa88}
Ge, Z.; Marsden, J.E.
\newblock {L}ie-{P}oisson {H}amilton-{J}acobi Theory and {L}ie-{P}oisson
  Integrators.
\newblock {\em Phys.\ Lett.\ A} {\bf 1988}, {\em 133},~134--139.

\bibitem{Sanz92}
Sanz-Serna, J.M.
\newblock The Numerical Integration of {H}amiltonian Systems.
\newblock  Computational Ordinary Differential Equations; Cash, J.R.; Gladwell,
  I., Eds.; Clarendon Press: Oxford,  1992; pp. 437--449.

\bibitem{HaLW02}
Hairer, E.; Lubich, C.; Wanner, G.
\newblock {\em Geometric Numerical Integration}, 2nd ed.; Vol.~31, {\em
  Springer Series in Computational Mathematics}, Springer: Berlin,  2006.
{
\bibitem{BC16}
Blanes, S.; Casas, F.
\newblock {\em A Concise Introduction to Geometric Numerical Integration}, 
CRC Press, Taylor \& Francis Group: Boca Raton, 2016.

\bibitem{HaLW03}
Hairer, E.; Lubich, C.; Wanner, G.
\newblock Geometric numerical integration
illustrated by the St\"ormer--Verlet method. 
\newblock {\em Acta Num.} {\bf 2003}, pp. 399--450.
}


\end{thebibliography}

\documentclass[preprints,commentary,accept,moreauthors,fleqn,pdftex]{Definitions/mdpi} 

\firstpage{1} 
\pubvolume{xx}
\issuenum{1}
\articlenumber{5}
\pubyear{2019}
\copyrightyear{2019}
\history{ }

\pdfoutput=1

\usepackage{latexsym}

\usepackage{amssymb}
\usepackage{amsfonts}

\numberwithin{equation}{section}
\numberwithin{theorem}{section}
\numberwithin{definition}{section}
\numberwithin{property}{section}

\newcommand{\dis}{\displaystyle}

\newcommand{\ba}{\begin{array}}
\newcommand{\ea}{\end{array}}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}

\newcommand{\ods}{\par \vspace{0.2cm} \par}

\newenvironment{my-itemize}{
  \begin{list}{$\bullet$}{
  \setlength{\itemsep}{10pt}
  \setlength{\parskip}{5pt}
  \setlength{\parsep}{15pt}
  \setlength{\leftmargin}{40pt}
  \setlength{\rightmargin}{0pt}}
}{
  \end{list}
}

\newcommand{\bi}{\begin{my-itemize}}
\newcommand{\ei}{\end{my-itemize}}

\Title{{On the famous unpublished } preprint \\
 ``Methods of integration which preserve the contact transformation property
 of the Hamilton equations'' \\ 
by Ren\'e De Vogelaere}

\Author{Robert D. Skeel $^{1}$, \ \    Jan L. Cie\'sli\'nski $^{2}$}

\AuthorNames{Robert D. Skeel and Jan L.\ Cie\'sli\'nski}

\address{%
$^{1}$ \quad School of Mathematical and Statistical Sciences,
Arizona State University, Tempe, AZ 85287, USA; rskeel@asu.edu \\ 
$^{2}$ \quad Uniwersytet w Bia\l ymstoku, Wydzia{\l} Fizyki, ul.\ Cio\l kowskiego 1L,  15-245 Bia\l ystok, Poland;  j.cieslinski@uwb.edu.pl}

\abstract{The well known 1956 unpublished report by De Vogelaere, motivated by the physics of particle accelerators, 
is the first demonstration of the existence of numerical integrators (now called symplectic integrators)  
that preserve a basic property of Hamiltonian systems. The purpose of this short note is to introduce Ren\'e De Vogelaere,
list his pioneering accomplishments,
and relate them to other early work on symplectic integrators. The preprint itself, typset in LaTeX, is attached as an appendix. }

\keyword{De Vogelaere; symplectic integrator; Hamiltonian systems}

\MSC{01A60; 65P10}

\begin{document}



Ren\'e De Vogelaere was born in Etterbeck, Belgium, on May 18, 1926, and died  in Berkeley, USA,  on December 14,  1991. He was a student of Georges Lema\^{i}tre  at the Catholic University of Louvain in Belgium,  receiving a Ph.D. in mathematics in 1948. 
He taught first at Laval University in Quebec,
then at Notre Dame University in Indiana,
where he wrote this well known and farsighted manuscript,
and finally, { since 1958,} at the University of California, Berkeley~\cite{CCLP92}.

The problem of numerically integrating Hamiltonian systems
on very long time intervals was brought to De Vogelaere's attention
by computational physicist J. Snyder.%
\footnote{Snyder
later headed the Computer Science Department at the University of Illinois.}
It was in connection with the design of a particle accelerator
to be built in the Midwestern United States.\footnote{
The project was later cancelled, soon after L.B. Johnson became president.}
De Vogelaere thought that if the numerical integrator
possessed the same qualitative properties as the Hamiltonian system,
then the effect of discretization error would be no worse than
the effect of errors in the model,
a kind of thinking associated with backward error analysis.
The resulting 1956 Notre Dame report~\cite{DeVo56}
produces a number of positive results on numerical symplectic integration.
It shows that the two variants of the method
that later became known as the symplectic Euler method
satisfy the property of being symplectic.
The report characterizes this method as being ``known'',
for the case of a {\em separable} Hamiltonian system.
The report also generalizes the method to a nonseparable system.
By composing the two variants of the nonseparable symplectic Euler method,
De Vogelaere obtains second order schemes for nonseparable systems.
These generalize each of the two variants
of the St\"ormer-Verlet/leapfrog method.
The extension of staggered schemes to nonseparable systems
is unique among those researchers who independently
discover the existence of symplectic numerical integrators.

There is no evidence of any further work on symplectic integration
until the 1983 article by R.D. Ruth~\cite{Ruth83},
who again is motivated by the dynamics of particle accelerators
but unaware of the unpublished report of De Vogelaere.
Using generating functions, Ruth shows that the leapfrog method is symplectic
and derives a novel third order symplectic method,
which employs a Jacobian matrix.
These results are for separable Hamiltonians.

Later that year and again unaware of prior work~\cite{ChSc90},
physicist P.J. Channell writes an internal report~\cite{Chan83}
showing the possibility of doing symplectic integration,
again using generating functions.

A year later, mathematician K.~Feng~\cite{Feng85,Feng86}
independently discovers symplectic numerical integrators,
in particular, the implicit Gauss-Legendre Runge-Kutta methods
and the (staggered) St\"ormer-Verlet method for separable Hamiltonians,
and illustrates the use of generating functions.
Publication was in a mathematics journal that was just three years old,
and the work initially received little attention.

However, by 1988,
a number of mathematical researchers had became
aware of and interested in the development
of symplectic (or canonical) integrators.
In particular, the seminal 1988 paper of Ge \& Marsden~\cite{GeMa88}
 acknowledges De Vogelaere, Channell, and Feng as developers
of the first symplectic integrators. { It is significant that in this paper  Marsden acknowledges also personal interactions with Ren\'e De Vogelaere, in fact they both were at Berkeley at that time.} 
Awareness of symplecticness
by the numerical ordinary differential equation community
became more widespread after the invited paper by J.M. Sanz-Serna
at the IMA Conference on Computational Differential Equations in 1989.
The published version~\cite{Sanz92} cites also the work of Ruth.
For a comprehensive account of subsequent developments,
consult Ref.~\cite{HaLW02}.

The report ``Methods of integration which preserve the contact transformation property  of the Hamilton equations'' \cite{DeVo56}, written in 1956 and still being an unpublished preprint, can be considered as a pioneering work in the field presently known as  geometric numerical integration, far ahead of its time.  {\it  ``This is a marvellous paper, short, clear, elegant, written in one week, submitted for publication -- and never published'' } (E. Hairer, C. Lubich and G. Wanner, see \cite{HaLW03}, p. 406). 

The content of the preprint can be easily summarized in few words, using modern terminology. Ren\'e De Vogelaere has shown that the methods presently known as symplectic Euler methods and St\"ormer-Verlet (or leap-frog) method are symplectic. 

Three cases are considered, in order of increasing complications:
\bi
   \item separable: \  $\ddot x = f (x)$
\item  non-separable: \  $\dot x = f (x,y)$  , \quad $\dot y = g (x,y)$  , \quad  $\displaystyle \frac{\partial f}{\partial y} + \frac{\partial g}{\partial x} = 0$  , 
\item multi-dimensional: \ $\displaystyle   \frac{d q_i}{d t} = \frac{\partial H}{\partial p_i}$  , \quad 
   $\displaystyle  \frac{d p_i}{d t} = - \frac{\partial H}{\partial q_i}$  , \quad 
   $(i = 1,\ldots,n)$  .
 \ei
In the first case notation $\ddot x = g (x)$ would be more convenient, because then this case be obtained from the second one as a special case $f (x,y)=y$, $g(x,y) = g (x)$.   

The symplectic Euler methods (Eqs. (3.1) and (3.3) in \cite{DeVo56}) are the only ones referered to as ``known methods'' by De Vogelaere. He extends the method (3.1) on the second and third case (see Eqs. (5.1) and (7.3) in \cite{DeVo56}) showing that they are symplectic of order 1. The extension of Eq. (3.3) on the second and third case is left as an obvious exercise. Hairer, Lubich and Wanner present these results as Theorem 3.3,  attributing it explicitly to Ren\'e De Vogelaere, see \cite{HaLW02}, p. 189. The role of De Vogelaere in the development of the symplectic Euler methods is discussed also by Blanes and Casas \cite{BC16}, p. 22--23.

Another family of symplectic integrators, leap-frog methods, has even longer history \cite{HaLW03}. In the separable (multidimensional) case they began to be widely used in molecular dynamics only after Verlet's publication in 1967, however they were rediscovered many times before by astronomers and physicists (e.g., by Delambre 1792, Encke 1860 and St\"ormer 1907) and also by\ldots De Vogelaere \cite{DeVo56}. Loup Verlet himself found that this method was in fact used by Isaac Newton (in {\em Principia}, 1687) to prove Kepler's second law \cite{HaLW03}. The surprising effectiveness of this method is due to its symplecticity and this propoerty has been first noticed and proved by De Vogelaere in \cite{DeVo56}. His Eqs. (4.1), (4.3),(6.1) and (7.4) represent, in fact, the leap-frog method:  
\[  \begin{array}{l}
\displaystyle 
\pmb{q}_{1/2} = \pmb{q}_0 + \frac{1}{2} h  \ \frac{\partial H}{\partial \pmb{p}} 
(\pmb{q}_{1/2}, \pmb{p}_0) \ , \qquad 
\pmb{p}_{1/2} = \pmb{p}_0 - \frac{1}{2} h \ \frac{\partial H}{\partial \pmb{q}} (\pmb{q}_{1/2}, \pmb{p}_0)\ , \\[4ex]
\displaystyle
\pmb{q}_1 = \pmb{q}_{1/2} + \frac{1}{2} h  \ \frac{\partial H}{\partial \pmb{p}} 
(\pmb{q}_{1/2}, \pmb{p}_1) \ , \qquad 
\pmb{p}_1 = \pmb{p}_{1/2} - \frac{1}{2} h \ \frac{\partial H}{\partial \pmb{q}} (\pmb{q}_{1/2}, \pmb{p}_1)\ .
\end{array} \]
As already suggested this is the first known construction of the leap-frog method in the case of general non-separable canonical Hamiltonian systems. 

The title of the preprint is a little bit misleading because De Vogelaere considers only transformations in the phase space $(q,p)$ treating time as a parameter. Now, such transformations are rather called symplectic (or canonical). Contact transformations,  usually related to the extended phase space (including the time variable), are more general.

The problem of extending symplectic integrators on non-autonomous (time-dependent) Hamiltonian systems is much more difficult and still only particular 
results are available. It seems that in this aspect the way of introducing time dependence of integrators presented in the famous preprint is not so fruitful as the symplecticity of the integrators in the autonomous case. 

Ren\'e De Vogelaere wrote his report in just one week. James Snyder brought this problem to his attention at the 
MURA\footnote{Midwestern Universities Research Association, a collaboration between 15 universities with the goal of designing and building a particle accelerator.} meeting of April 14, 1956 and the preprint appeared on April 21, 1956.   
Taking into account the surprisingly short time of writing this report, the number of mistakes seems to be very few. Actually, we noticed only one place with some computational errors (not very essential and without any influence on the conclusions of the paper): coefficents by several terms of the operator $\Phi$, appearing in Eq.~(6.2) in section 6, are not correct. It should read as follows:
\[
\begin{array}{l}  \displaystyle
\Phi = - \frac{1}{24} f^2 \frac{\partial^2}{\partial x^2} + \frac{1}{12} g^2 \frac{\partial^2}{\partial y^2} + \left( \frac{\alpha^2 - \alpha}{2} + \frac{1}{12} \right) \frac{\partial^2}{\partial t^2}     + \left(  \frac{1}{6} - \frac{\alpha}{2}  \right) g  \frac{\partial^2}{\partial y \, \partial t} 
\\[3ex] \dis \qquad  - \frac{1}{12} f \frac{\partial^2}{\partial x \, \partial t}  -  \frac{1}{12} f g \frac{\partial^2}{\partial x \, \partial y} 
 + \frac{1}{12} \frac{\partial g}{\partial t} \frac{\partial }{\partial y} + \left( \frac{\alpha}{2} - \frac{1}{6} \right) \frac{\partial f}{\partial t} \frac{\partial }{\partial x} 
 \\[4ex] \displaystyle \qquad
 + \frac{1}{12} f  \frac{\partial g}{\partial x} \frac{\partial }{\partial y} 
 + \frac{1}{12} f  \frac{\partial f}{\partial x} \frac{\partial }{\partial x} + \frac{1}{12} g  \frac{\partial g}{\partial y} \frac{\partial }{\partial y} - \frac{1}{6} g \frac{\partial f}{\partial y}  \frac{\partial }{\partial x} \ . 
\end{array} 
\]

The De Vogelaere's preprint \cite{DeVo56}, typeset in LaTeX for Readers' convenience (the original report is written on a  mechanical typewriter and is hardly readable in some places), is  available as a supplementary material  (appendix)  to our short article.  We tried to preserve the original form of the manuscript, except correcting few obvious misprints and taking also into account the  handwritten corrections made on the preprint, probably by the author  himself. 

\ods\ods



\acknowledgments{We thank Charles DeVogelaere, PhD, and the family of Ren\'e DeVogelaere for granting permission to publish the preprint \cite{DeVo56}. Many thanks are due to the University of Notre Dame and, especially, to J.~Parker Ladwig,  for assistance and help.}

\ods\ods



\reftitle{References}

\end{document}